if lynn can type a page in p minutes, what piece of the page can she do in 5 minutes: if lynn can type a page in p minutes, what piece of the page can she do in 5 minutes
Answer by oscargut(667)

1 page in p minutes
x pages in 5 minutes
then x=1*5/p=5/p
Answer: 5/p pages